# Local Conditional Controlling for Text-to-Image Diffusion Models

### v1

@@ 3.5 Focused Token Response
After updating the noised latent, while the image content closely aligns with the prompt, objects still tend to be generated near the local control region. This challenge is often linked to the overlapping contribution areas of corresponding token embeddings. To enhance object distinction, inspired by[36], we utilize cross attention maps to mitigate weak response regions. Specifically, we reduce the attention scores for patchs in the attention map that do not demonstrate the strongest response in the token dimension. This technology encourages the reduction of responses from less dominant tokens, thereby enforce the model to selectively respond to the most relevant tokens.
$\mathcal{A}^{t}=O(\arg\max(\mathcal{A}^{t}),\gamma)\cdot\mathcal{A}^{t}$ (6)
Here, $\arg\max$ identifies the indices of the highest responses across token dimension, while $O$ represents a suppression operation that applies a scaling factor $\gamma$ to the non-max attention scores. This process reduces the impact of irrelevant pixels in the cross attention map for the current token.
@@
3.6 Feature Mask Constraint
During the denoising process with local control, there is a notable difference in the amount of information between inside and outside of the local control region. The local control carries only localized control information, leaving the rest of the areas with blank control information. As a consequence, the latent distribution resulting from the injection of local control conditions into the U-Net differs from that obtained under global control conditions, potentially affecting the overall image quality.
To address this issue, we apply feature mask constraint to the output of ControlNet. This design specifically directs the U-Net’s focus to the local control region when adding control. It thus minimizes the disruptive effect of irrelevant blank control information in non-control region. In practical terms, we extract the local feature maps from the output of ControlNet block. Subsequently, these local condition features were added in the Unet. Suppose $\mathcal{F}(\cdot)$ is a neural block in U-Net where $x$ is the input feature map. Separately, $\mathcal{G}(\cdot)$ is a neural block in ControlNet, and the conditional input feature $c_{f}$ is derived from $\mathcal{C}_{f}$. The resultant feature map $y$ includes added control conditions.
$y=\mathcal{F}(x)+\mathbf{M}\cdot\mathcal{G}(x,c_{f})$ (7)
After applying the feature mask constraint, the output from the ControlNet is selectively integrated into the U-Net. The local control condition is achieved through a multiplication operation involving the feature mask $\mathbf{M}$ and the conditioned output from the ControlNet block, $\mathcal{G}(x,c_{f})$. This combination is then added to the original feature map $\mathcal{F}(x)$ from the U-Net. By explicitly preventing the U-Net from focusing on blank control information, this method effectively mitigates image quality issues arising from disparities in control information.
@@ 4.1 Dataset and Evaluation
Dataset: We utilized the COCO[19] validation set, which comprises 80 object categories. Each image in this set is associated with multiple captions; for our evaluation, we randomly selected one caption per image, resulting in a dataset of 5k generated images. Additionally, we create a dataset inspired by the Attend-and-Excite [2], named Attend-Condition. This dataset encompasses 11 object and 11 animal categories, with text prompts typically structured as ‘a [object] and a [animal]’. We selected one representative image for each concept to serve as control information for local control, generating two images per prompt, totaling 968 validation data.
@@
Evaluation Metrics: For assessing image quality, we used the Fréchet Inception Distance (FID)[32], while the average CLIP[23] score between text prompts and corresponding images measured text-image similarity. The primary focus of our evaluation is on whether the generated images align with the objects mentioned in the captions. Utilizing the descriptive capabilities of BLIP-2[18], we generate captions for the images produced under the local control paradigm. Subsequently, we calculate the CLIP text-text similarity between these generated captions and the original captions, providing a comprehensive measure of content alignment with the text prompt.
Dataset Condition FID $\downarrow$ CLIP Score $\uparrow$ CLIP T2T $\uparrow$
ControlNet T2I-Adapter Ours ControlNet T2I-Adapter Ours ControlNet T2I-Adapter Ours
COCO2017 Canny 23.65 28.86 21.38 0.283 0.295 0.303 0.800 0.777 0.795
Depth 32.08 26.57 22.01 0.290 0.298 0.307 0.763 0.788 0.804
Sketch 25.96 34.06 22.33 0.303 0.292 0.307 0.799 0.771 0.806
Attend-Condition Canny - - - 0.271 0.254 0.287 0.750 0.694 0.802
Depth - - - 0.251 0.261 0.283 0.713 0.714 0.806
Sketch - - - 0.264 0.256 0.286 0.741 0.685 0.801
Table 1: Quantitative comparisons to other controllable methods: ControlNet [44] and T2I-Adapter [21]. We do not provide FID on the Attend-Condition dataset due to the small scale of its original images. Our method achieves the best performance under most metrics.
Figure 6: Comparisons to different controllable diffusion models on local control.
4.2 Visualization Results
Figure 4 demonstrates the practical utility of our method in achieving local control. It ensures that the generation process remains attentive to local content control, while simultaneously allowing the rest of the regions to generate content accurately in accordance with the prompt. Compared to images generated by ControlNet[44], our approach effectively reduces the issue of local control dominance.
Figure 5 further showcases the flexibility of our approach. It presents images generated under various conditions, showing the method’s ability to seamlessly integrate with the existing ControlNet framework. This flexibility highlights our approach’s capability to synthesize images accurately under a wide range of control conditions.
Figure 7: Ablation Studies. We use ControlNet[44] as our baseline. In the fourth column, we add Object Regeneration (OR) to baseline method. The fifth column builds upon the previous one by incorporating Focused Token Response (FTR). Finally, the sixth column extends the previous setup by adding Feature Mask Constraint (FMC).
4.3 Comparisons
We conducted a comparative analysis of our method against ControlNet[44] and T2I-Adapter[21], using local control condition as the input condition. This involved evaluating the methods under various conditions such as canny, depth, and sketch on the COCO and Attend-Condition datasets. The results, encompassing metrics FID, CLIP-score and CLIP text-text similarity (CLIP T2T), are detailed in Table 1. Due to the small number of original images in the Attend-Condition, we do not calculate its FID, as such a small sample size leads to unreliable FID. Our method outperformed the aforementioned control methods in most conditions under the paradigm of local control.
Figure 6 illustrates these differences, particularly in scenarios involving multiple objects in prompts. In the first example, ControlNet and T2I-Adapter tended to merge the concepts of a building and an airplane into a single object within the local control region. In contrast, our method generate a building in line with the local control conditions while accurately generating the plane in the non-local control area. Another striking example is seen in the third row on the left, demonstrates ControlNet’s output where it mistakenly generates part of a rabbit’s shape on a turtle. Our method, however, showcased its advanced capability to distinctly recognize and generate separate objects. This allows for greater flexibility in producing user-specified images within the local control.
@@ 4.4 Ablation Studies
In our ablation study, we utilize the canny condition on the COCO dataset to evaluate the effectiveness of our method’s key components. Specifically, we focus on three compon ents: 1) Object Regeneration, updates the noised latent to assist object generation in non-local regions; 2) Focused Token Response, designed to refine attention response; and 3) Feature Mask Constraint, a mechanism to constraint features of the ControlNet output. The results are listed in Table 2. It reveals that updating noised latent and refine attention response effectively reduce the influence of local control dominance in non-local regions. This enables more accurate image generation in accordance with the given prompts, as evidenced by a high CLIP T2T score of 0.806. Since non-local regions have blank control information, the gap in control conditions between inside and outside the local control regions impacts image quality. Our study shows that incorporating feature mask constraint significantly mitigates this issue. This integration resulted in notable improvements in both CLIP and FID scores, with CLIP scores rising to 0.303 and FID scores decreasing to 21.38, indicating enhanced overall image quality.
OR FTR FMC FID CLIP Score CLIP T2T
- - - 23.65 0.283 0.750
$\surd$ - - 24.48 0.299 0.786
$\surd$ $\surd$ - 22.90 0.298 0.806
$\surd$ $\surd$ $\surd$ 21.38 0.303 0.795
Table 2: Ablation studies on the proposed components, including Object Regeneration (OR), Focused Token Response (FTR), and Feature Mask Constraint (FMC). All components are crucial.
Figure 7 visualizes the impact of progressively integrating our designs on image generation. In the first row, using only the local control condition results in generating a train that matches its spatial structure, but the car is missing. Adding Attention Contrast Enhancement improves this, allowing the car mentioned in the prompt to appear, although some overlap and sticking issues are still evident. When Focused Token Response is incorporated, the image correctly produces all concepts from the prompt. Moreover, the introduction of the feature mask constraint significantly enhances image quality. Similar improvements are observed in other examples, confirming the effectiveness of our method’s components.
5 Limitations
The limitation pertains to the semantic connection between inside and outside of the local control region. This can cause problems with how objects are sized and placed, making them look out of proportion. Future work could address this by establishing a more comprehensive connection between the inside and outside of the local control regions.
@@ 6 Conclusions
In this paper, we introduce local control as a new paradigm for image controllable generation, focusing on user-defined control of specific image areas. To mitigate local control dominance, we propose a training-free method that integrates smoothly with existing control models. This approach optimizes noised latents and cross-attention maps, enhancing object generation in non-local regions. We demonstrate a variety of results under the local control condition, successfully generating high-quality images faithful to the text prompts.
Supplementary Material
Considering the space limitation of the main text, we provided more results and discussion in this supplementary material, which is organized as follows:
  • $\bullet$
    Section 7: implementation details.
  • $\bullet$
    Section 8: more ablation studies
    • discussion on Control Concept Matching.
    • selection of the $\gamma$ value in Focused Token Response
  • $\bullet$
    Section 9: more visualization results.
    • visualization results for COCO and Attend-Condition.
    • visualization results for different condition and prompt.
  • $\bullet$
    Section 10: the necessity of the local control.
7 Implementation Details
Experiments.
All experimental results are obtained using the official Stable Diffusion [27] v1.5 text-to-image model and ControlNet [44]. The number of denoising steps $T$ is set as 50 while configuring the classifier-free guidance to 7.5. $\beta$ is set as 0.8 in Control Concept Matching , which means that $\beta T$ is 40. A Gaussian kernel $G(\cdot)$ with a size of $3\times 3$ and a standard deviation $\sigma=0.5$ is used to smooth the attention map. To reduce the attention scores of non-maximum values, $\gamma$ is set to 0.85. We use spaCy [13] which is a library for advanced NLP, to extract noun tokens based on the part of speech of each word in the caption. All experiments are conducted on a NVIDIA A100 GPU.
Datasets.
As discussed in Section 4.1 of the main paper, we define Attend-Condition to further validate our method’s effectiveness in generating multi-object scenes under the paradigm of local control. In Table  3, we list the animals and objects inspired by [2].
Animals
cat, dog, bird, bear, horse, elephant, monkey,
frog, turtle, rabbit, mouse
Objects
glasses, crown, suitcase, chair, balloon,
car, bowl, bench, clock, apple
Table 3: Datasets. We list the animals and objects used in Attend-Condition dataset.
In the COCO [19] validation dataset, we use the largest segment region in each image as the local spatial information for inference. Specifically, we generate a binary mask from the image’s largest segment region. This mask is then combined with the global control condition, thereby defining the local control condition. The text prompt is randomly selected from one of the five captions corresponding to the current image. Then the local control condition and text prompt are fed into our pipeline.
8 More Ablation Studies
Control Concept Matching
The figure 8 shows the necessity of selecting concepts which match the local control region. The middle images correspond to manually specified $\text{C}^{t}_{\text{control}}$ of “cat” and “dog” , respectively. It can be observed that the content generated in the local region aligns with the manually specified objects. However, this naive method leads to the forced generation of objects that should not be in local region, affecting the model’s ability to align with the control condition. The results for both cat and dog generation are bad. The right image demonstrates the superiority of Control Concept Matching. It adaptively matches the concept during denoising process. More examples are shown in the Figure 9. Furthermore, Figure 10 shows the impact of different $\beta$ values on image quality.
Figure 8: Comparison of $\text{C}^{t}_{\text{control}}$ obtained by manually specifying and Control Concept Matching. Our method does not require manually specifying $\text{C}^{t}_{\text{control}}$ and produces higher quality images.
Focused Token Response
In Figure 11, we demonstrate the impact of different values of $\gamma$ on the synthesized image. When $\gamma$ is 1, it is difficult to generate images with both “cat” and “dog” , and alignment with the local condition is not achieved. When $\gamma$ is set to 0.85, the generated images meet expectations. As $\gamma$ decreases from 0.7 to lower values, the overall quality of the images significantly deteriorates.
9 More Visualization Results
We present visualization results under local control on both COCO validation and Attend-Condition datasets. Figure 14 displays Attend-Condition examples, showcasing the ability to align with captions and local conditions across various concept combinations. In Figure 13, selected original images of varying sizes from the COCO validation set and their corresponding generated images are displayed. Our method controls the generation of specific image areas based on user-defined local conditions, with the rest of the image following the original text prompt. Moreover, in the last two rows of Figure 13, even when the local control condition almost covers the entire image, our method still correctly generates images that comply with both the prompt and control condition. This demonstrates that our approach can generalize to global control tasks. Figure 12 provides additional visualization results applied under various conditions and prompts.
10 The Necessity of the Local Control
We emphasize the necessity of the local control paradigm. In Figure 15, when global control information at the image-level is input, the model can generate good content for the main subject as dictated by the control information. However, other objects described in the prompt depend on the control information outside the main subject’s control area. When the caption does not match well with the remaining control information, it can result in failure cases or produce somewhat strange content.
Figure 9: The first row represents manually specifying $\text{C}^{t}_{\text{control}}$ as “dog”, while the second row employs the proposed Control Concept Matching method. Our method generates concepts in the local region. At the same time, it correctly creates other parts of the image according to the prompt. It does not generate the same concept twice or create extra concepts.
Figure 10: Different $\beta$ values affect image quality. Our focus is on stabilizing the selection of $\text{C}^{t}_{\text{control}}$. With a higher $\beta$, emphasizing local region objects too early leads to poorer quality in these areas. Conversely, a lower $\beta$ means the objects assigned within the local region during the denoising process change, resulting in unstable content generation in the images. We have chosen 0.8 for our method.
Figure 11: The impact of different values of $\gamma$ on the synthesized image. At $\gamma=0.85$, image quality is optimal, and it declines below 0.7 and struggles with combination object generation at $\gamma=1$.
Figure 12: More visualization results. Our method achieves excellent results under a variety of conditions and prompts.
Figure 13: Synthetic samples with local conditions on COCO validation set. Our method accurately generates images under different control regions. Additionally, when the control covers almost the entire image (in the last two rows), our approach is still effective.
Figure 14: Synthetic samples with local conditions on Attend-Condition. The images showcase our method’s ability to align with captions and local conditions across diverse concept combinations.
Figure 15: Under global conditions, “cat” can be generated normally, while “toy vehicle” can only be generated based on background control conditions. The final generated result is unsatisfactory and does not meet the user’s needs. In the local control paradigm, users can specify “cat” to use the control condition, while the remaining areas are generated correctly based on the prompt.

### v3

@@ 3.4 Focused Token Response
After updating the noised latent, while the image content closely aligns with the text prompt, non-local objects still tend to be generated repeatedly, or near the local control region. This challenge is often linked to the overlapping contribution regions of corresponding token embeddings. Inspired by(Tunanyan et al. 2023), we utilize cross attention maps to mitigate weak response regions. Specifically, we reduce the attention scores for patchs in the attention map that do not demonstrate the strongest response in the token dimension. This technology encourages the reduction of responses from weaker tokens, thereby enforce the model to selectively respond to the most relevant tokens.
$\mathbf{A}^{t}=O(\arg\max(\mathbf{A}^{t}),\gamma)\cdot\mathbf{A}^{t},$ (8)
$\arg\max$ identifies the indices of the highest responses across token dimension, while $O$ represents a suppression operation that applies a scaling factor $\gamma$ to the non-max attention scores. This technique minimizes the influence of irrelevant pixels in the cross-attention map, enhances specific token contributions, reduces overlap and duplication. Collaborating with RDLoss, it aids the model in generating across both local and non-local regions.
3.5 Feature Mask Constraint
During the denoising process, there is a notable difference in the amount of information between inside and outside of the local control region. The local control carries only localized control information, leaving the rest of the regions with blank control information. As a consequence, the latent distribution resulting from the injection of local control conditions into the UNet differs from that obtained under global control conditions, potentially affecting the image quality.
To address this issue, we apply feature mask constraint to the output of ControlNet. This design specifically directs the UNet’s focus to the local control region when adding control. It thus minimizes the disruptive effect of irrelevant blank control information in non-control region. In practical terms, we extract the local feature maps from the output of ControlNet block. Subsequently, these local condition features are added in the UNet. Assume $\mathcal{F}(\cdot;\theta)$ represents a neural block within the UNet architecture, and $x_{t}$ is the input feature map of the block during the $t$-th denoising step. Separately, $\mathcal{G}(\cdot;\cdot)$ represents a neural block in ControlNet, which takes the spatial conditional input feature $c_{f}$ and $x_{t}$. The fused feature output from each block is denoted as $y_{t}$:
$y_{t}=\mathcal{F}(x_{t};\theta)+\mathbf{M}\cdot\mathcal{G}(x_{t};c_{f}).$ (9)
This technique is applied in the computation process of the RDLoss. After applying the feature mask constraint, the output from the ControlNet is selectively integrated into the UNet. The local control condition is achieved through a multiplication operation involving the feature mask $\mathbf{M}$ and the conditioned output from the ControlNet block $\mathcal{G}(x_{t};c_{f})$. This combination is then added to the original feature map $\mathcal{F}(x_{t};\theta)$ from the UNet. This method improves image quality by explicitly preventing the UNet from focusing on irrelevant control information.
@@ 3.5 Feature Mask Constraint
Figure 5: Comparisons with several baselines under varying conditions. Our method, along with T2I-Adapter and ControlNet, uses local conditions as input. Conversely, Noise-Mask, Feature-Mask, and Inpainting use global conditions as inputs. All these methods use the same control mask. Compared with these baselines, our method can synthesize high-quality images in local control, especially in terms of text alignment.
Dataset metrics ControlNet T2I Noise-Mask Feature-Mask Inpainting Ours
COCO2017 FID $\downarrow$ 27.87 $\pm$ 4.22 30.32 $\pm$ 3.75 22.72 $\pm$ 0.61 21.93 $\pm$ 0.64 25.72 $\pm$ 1.22 21.86 $\pm$ 0.48
CLIP Score $\uparrow$ 0.293 $\pm$ 0.010 0.295 $\pm$ 0.03 0.287 $\pm$ 0.003 0.290 $\pm$ 0.002 0.288 $\pm$ 0.003 0.305 $\pm$ 0.02
CLIP T2T $\uparrow$ 0.782 $\pm$ 0.019 0.780 $\pm$ 0.009 0.789 $\pm$ 0.006 0.779 $\pm$ 0.014 0.785 $\pm$ 0.006 0.801 $\pm$ 0.006
IOU $\uparrow$ 0.55 $\pm$ 0.07 0.44 $\pm$ 0.06 0.41 $\pm$ 0.02 0.37 $\pm$ 0.03 0.53 $\pm$ 0.07 0.51 $\pm$ 0.10
LPIPS×${}_{{}^{\times 10^{2}}}$downarrow 13.15 $\pm$ 1.19 14.00 $\pm$ 1.06 14.16 $\pm$ 1.12 15.01 $\pm$ 1.33 13.25 $\pm$ 1.51 13.04 $\pm$ 1.36
Attend-Condition FID $\downarrow$ - - - - - -
CLIP Score $\uparrow$ 0.261 $\pm$ 0.010 0.258 $\pm$ 0.004 0.279 $\pm$ 0.003 0.278 $\pm$ 0.003 0.273 $\pm$ 0.03 0.285 $\pm$ 0.02
CLIP T2T $\uparrow$ 0.732 $\pm$ 0.018 0.700 $\pm$ 0.015 0.782 $\pm$ 0.006 0.784 $\pm$ 0.019 0.753 $\pm$ 0.009 0.804 $\pm$ 0.002
IOU $\uparrow$ 0.74 $\pm$ 0.04 0.70 $\pm$ 0.02 0.66 $\pm$ 0.02 0.64 $\pm$ 0.04 0.72 $\pm$ 0.04 0.70 $\pm$ 0.05
LPIPS×${}_{{}^{\times 10^{2}}}$downarrow 12.67 $\pm$ 1.03 13.31 $\pm$ 1.15 13.25 $\pm$ 1.21 13.30 $\pm$ 1.27 12.97 $\pm$ 1.06 12.57 $\pm$ 1.08
Table 1: Quantitative comparisons to baselines. We do not provide FID on the Attend-Condition dataset due to the small scale of its original images. Our method can maintain good alignment with the provided spatial condition while achieving the best image quality and text alignment.
Figure 6: Comparisons with baselines on Attend-Condition.
@@ 4.1 Dataset and Evaluation
Dataset: We utilized the COCO(Lin et al. 2014) validation set, which comprises 80 object categories. Each image in this set is associated with multiple captions, we randomly selected one caption per image, resulting in a dataset of 5k generated images. Additionally, we create a dataset inspired by (Chefer et al. 2023), named Attend-Condition. This dataset encompasses 11 object and 11 animal categories, with text prompts typically structured as ‘a [object] and a [animal]’. We collect images of each concept from the internet and randomly select those with a larger subject area to serve as representative images for local control.
@@
Evaluation Metrics: For assessing image quality, we used the Fréchet Inception Distance (FID)(Seitzer 2020), while the average CLIP(Radford et al. 2021c) score between text prompts and corresponding images measured text-image similarity. Additionally, the primary focus of our evaluation is on whether the generated images align with the objects mentioned in the captions. Utilizing the descriptive capabilities of BLIP-2(Li et al. 2023a), we generate captions for the images produced under the local control paradigm. Subsequently, we calculate the CLIP text-text similarity between these generated captions and the original captions, providing a comprehensive measure of content alignment with the text prompt. To evaluate the alignment with the provided spatial condition, we caculate the LPIPS(Zhang et al. 2018) metrics between original and generated images in localized regions. To evaluate if local object generation matches specified regions, we use the IOU metric. we adopt the segmentation-based evaluation approach used in ControlNet on the COCO dataset, which uses OneFormer(Jain et al. 2023) to detect the local segmentations then compute the IOU. Additionally, we use SAM(Kirillov et al. 2023) and DINO(Liu et al. 2023) to segment the Attend-Condition experimental results.
Baseline Settings We conducted a comparative analysis of our method against several baselines. We select two commonly used structure-controllable models, ControlNet(Zhang, Rao, and Agrawala 2023) and T2I-Adapter(Mou et al. 2023), as naive baselines by inputting the local control condition and text prompts into these models. Additionally, we design three competitive baselines specifically for the local control task, which do not require extra training and use global conditions as input. Specifically, 1)Noise-Mask: Inspired by (Nichol et al. 2021), we estimate the noise with adding the condition into ControlNet and the noise from vanilla T2I model, then combine them according to the mask. The combined noise is calculated as:
$\epsilon_{\theta}=\epsilon_{\theta}(\bm{z}_{t},\bm{t},\bm{c},\bm{c}_{\text{f}})\cdot\mathbf{M}+\epsilon_{\theta}(\bm{z}_{t},\bm{t},\bm{c})\cdot(1-\mathbf{M}).$ (10)
2)Feature-Mask: We implement the addition of our proposed “Feature Mask Constraint” component into ControlNet across all denoising steps, which applies local control mask to ControlNet features. 3)Inpainting: We use ControlNet to generate an image with the spatial condition and text prompt, then use inpainting method(Lugmayr et al. 2022) to regenerate the parts without spatial condition.
4.2 Comparison with Baselines
Figure 7: Ablation Studies: a) We use ControlNet(Zhang, Rao, and Agrawala 2023) as the vanilla baseline. b) Only RDLoss is included. c) Only the FMC is included, referred to as the feature-mask baseline. d) Our method without the FTR module. e) Our method without the FMC module. f) Our complete method.
Quantitative experiments. We evaluate the baselines and our method under various conditions such as canny, depth, and sketch on the COCO and Attend-Condition datasets. The result includes the following metrics: FID, CLIP-score, CLIP text-text similarity (CLIP T2T), IOU, and LPIPS, which are detailed in Table 1. Due to the small number of original images in the Attend-Condition, we do not calculate its FID, as such a small sample size leads to unreliable FID. In certain cases, the local control region is significantly smaller than the entire condition image, which results in many blank pixels. This leads to the naive methods generating artifacts and grayscale images, which result in high FID scores. In evaluations on Attend-Condition dataset specifically designed for multi-object generation, our method significantly outperforms other baselines in both CLIP Score and CLIP T2T. Additionally, we evaluate the alignment ability of baselines and our method. Noise-Mask and Feature-Mask methods align poorly with the given local spatial conditions. ControlNet achieves the best spatial alignment performance. The pipeline of the inpainting method ensures spatial alignment capabilities similar to that of ControlNet. Moreover, our method also achieves similar spatial alignment capabilities as ControlNet in local control task. Compared to these baselines, our method maintains good alignment with the given spatial conditions, while achieving the best image quality and text alignment.
Qualitative Results. Figure 5 illustrates these differences, particularly in scenarios involving multiple objects in prompts. In naive method, it can almost only generate objects in text prompt that best match the control region, while neglecting the generation of other objects. The other three baselines designed specifically for local control achieved good text-image alignment and image quality, but have significant drawbacks compared to our visualized results. Noise-Mask frequently generates untrustworthy pixels near the local region, and sometimes fails to generate mutilple objects. The results generated by the Feature-Mask approach align poorly with the given spatial conditions. The inpainting method sometimes fails to distinguish between the generated content within controlled and uncontrolled regions, leading to issues of repetition and omission in the generated images. For instance, an additional chair in the second row, and a turtle is missing in the third row. Compared to these baselines, our method generates images that most closely match the text prompts.
@@ 4.3 Ablation Studies
Main ablations. We utilize the canny condition on the COCO dataset to evaluate the effectiveness of our method’s key components. Specifically, we focus on three components: 1) Regional Discriminate Loss(RDLoss), updates the noised latent to assist object generation in non-local regions; 2) Focused Token Response(FTR), designed to refine attention response; and 3) Feature Mask Constraint(FMC), a mechanism to constraint features of the ControlNet output. Only the FMC is included, referred to as the feature-mask baseline. Additionally, since the FTR depends on RDLoss, there is no ablation study for it alone. The images are generated using the same seed. The results are listed in Table 2. It reveals that RDLoss reduces the influence of local control dominance in non-local regions. FTR and FMC can assist RDLoss in improving the quality of generated images. It can be foreseen that the updates to the latent have somewhat diminished the capability for image alignment, yet the extent of this impact remains within an acceptable threshold. Figure 7 visualizes the results of the ablation experiments.
[…]
- - - 23.65 0.283 0.750 0.62 11.96
$\surd$ - - 24.48 0.299 0.786 0.59 11.99
- - $\surd$ 21.83 0.292 0.791 0.40 13.68
$\surd$ - $\surd$ 21.91 0.301 0.802 0.57 11.71
$\surd$ $\surd$ - 22.90 0.298 0.806 0.59 11.79
$\surd$ $\surd$ $\surd$ 21.38 0.303 0.795 0.57 11.68
Table 2: Ablation studies of the proposed components under canny condition on the COCO dataset. Since the FTR depends on RDLoss, there is no ablation study for it alone.
Figure 8: (a) Control Concept Matching adaptively matches the concept, (b) Different $\beta$ values affect image quality.
Impact of Control Concept Matching. The Figure 8(a) shows the necessity of selecting concepts which match the local control region. It can be observed that the content generated in the local region aligns with the manually specified objects. However, this naive method leads to the forced generation of objects that should not be in local region, affecting the model’s ability to align with the control condition. The results for both cat and dog generation are bad. The right image shows the superiority of Control Concept Matching. It adaptively matches the concept during denoising process. Figure 8(b) shows the impact of different $\beta$ values on image quality. Our focus is on stabilizing the selection of $\text{C}^{t}_{\text{control}}$. A $\beta$ between 0.8 to 0.9 yields good results.
@@ 5 Conclusions
In this paper, we introduce local control as a new paradigm for image controllable generation, focusing on user-defined control of specific image regions. To mitigate local control dominance, we propose a training-free method that integrates smoothly with existing control models. This approach optimizes noised latents, enhancing object generation in non-local regions. We demonstrate a variety of results under the local control condition, successfully generating high-quality images faithful to the text prompts.